x-4y=2
-2x-5y=9
solve equations by substitution

Answer by ikleyn(52779)

.
x-4y=2
-2x-5y=9
solve equations by substitution
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

  x - 4y = 2,     (1)
-2x - 5y = 9.     (2)


From (1), express x = 2 + 4y and substitute it into (2). You will get a single equation for y:

-2*(2+4y) - 5y = 9.

Simplify and solve for y:

-4 - 8y - 5y = 9,

-13y = 9 + 4,

-13y = 13,

y = -1.

Then x = 2 + 4y = 2 + 4*(-1) = 2 - 4 = -2.

Answer. x = -2, y = -1.